# Group Distributionally Robust Dataset Distillation with Risk Minimization

### v2

@@ 3.1 Analysis
=$\mathbb{Q}=\{\mathcal{Q}_{i}\},\,\text{ with }Q_{i}=\text{supp}(\mathcal{Q}_{i}),\,\bigcup_{i=1}^{q}Q_{i}=D,\,\underline{s}\leq\lambda(Q_{i})\leq\bar{s}$
@@
The population distribution $\mathcal{D}$ has bounded support, i.e., $\lambda(D)<\infty$. The loss $F$ is Lipschitz continuously differentiable with respect to the first argument (the parameters) and continuous with respect to the second argument (the input features and labels).
@@
$\lim\sup\limits_{t\to\infty}(\text{supp}(\mathcal{Q})\cap\bar{\mathcal{T}}_{t})=\text{supp}(\mathcal{Q})$
@@
$\begin{array}[]{rl}\min\limits_{\mathcal{S}}\max\limits_{\mathcal{Q}\in\mathbb{Q}}&\mathbb{E}[F(\theta^{*},\mathcal{Q})]\\ \text{s.t. }&\theta^{*}\in\arg\min_{\theta}F(\theta,\mathcal{S}).\end{array}$ (3)
Notice that there are two layers of generalization: on the one hand, we are solving a problem robust with resepect to the choice of $\mathcal{Q}\in\mathbb{Q}$, and on the other hand, we are considering the population error rather than an empirical loss.
@@
that is ultimately used to steer the synthetic dataset, as the DD task is finding a dataset on whose associated $\theta$-minimizer also minimizes this quantity. In the analysis, we shall present:
(1) A justification for why our Algorithm approximately solves optimization problem (3).
(2) The discussion of the group robustness properties associated with the solution of (3).
To begin with, we rewrite (3) as:
$\begin{array}[]{rl}\min\limits_{\mathcal{S}}\max\limits_{\mathcal{Q}\in\mathbb{Q}}\max_{\mathcal{Q}^{\prime}\in\bar{\mathcal{Q}}}&\mathbb{E}[F(\theta^{*},\mathcal{Q}^{\prime})]\\ \text{s.t. }&\theta^{*}\in\arg\min_{\theta}F(\theta,\mathcal{S})\\ &\bar{\mathcal{Q}}=\left\{\mathcal{Q}^{\prime}:I(\mathcal{Q}^{\prime},\mathcal{S}\cup\mathcal{Q}_{N})\leq r\right\}\end{array}$ (5)
where $r>0$ is some bound and we replace the inner problem to be evaluated on the entire training dataset with a data-driven DRO. In the appendix we present several results in the literature that indicate how the DRO to an empirical risk minimization problem exhibits generalization guarantees and hence is a valid auxiliary criterion for minimizing (3).
@@ 3.2 Large Deviations and Solving the bilevel DRO
To justify the clustering and risk measure minimization algorithm as an appropriate procedure for solving (5), we apply some theoretical analysis on the relationship between DRO and Large Deviations Principles (LDP). LDPs (e.g. (Deuschel & Stroock, 2001)) define an exponential asymptotic decay of the measure of the tails of an empirical distribution with respect to a sample size.
@@
$\begin{array}[]{l}\lim\sup\limits_{N\to\infty}\frac{1}{N}\log\mathbb{P}_{\mathcal{Q}}^{\infty}\left(F(\theta^{*}(\mathcal{S}),\mathbb{P}_{\mathcal{Q}})>F(\theta^{*}(\mathcal{S}),\mathcal{S}\cup\mathcal{Q}_{N})\right)\\ \qquad\leq-r,\,\forall\mathcal{Q}\in\mathbb{Q}\end{array}$ (6)
for some $r>0$. This states that for all partitions $\mathcal{Q}$ of the population space partition $\mathbb{Q}$, there is an exponential asymptotic decay of the probability of one-sided sample error (i.e., samples from the population $\mathbb{P}_{\mathcal{Q}}$) relative to the computed loss on the synthetic and training data points $\mathcal{Q}_{N}$, as the number of training data points grows asymptotically.
@@
Algorithm (2) satisfies, under Assumptions 3.1 and 3.2, the LDP (7).
Next, we relate Algorithm (2) together with 3.1 to the DRO problem (5).
@@ 3.3 Large Deviations and Distributionally Robust Optimization
Applying Assumption 3.2, the LDP, and the continuity of $F$ with respect to the data, Assumption 3.1 (see also the proof of Lemma 1 as well as Example 3 in (Duchi & Namkoong, 2021)) we can bound the quantity:
$\mathbb{P}(F(\theta^{*},q)>F(\theta^{*},\mathcal{S}\cup\mathcal{Q}_{N}))$
@@
This implies that one can bound the objective value of the DRO problem of interest (5):
$\begin{array}[]{rl}\min\limits_{\mathcal{S}}\max\limits_{\mathcal{Q}\in\mathbb{Q}}\max_{\mathcal{Q}^{\prime}\in\bar{\mathcal{Q}}}&\mathbb{E}[F(\theta^{*},\mathcal{Q}^{\prime})]\\ \text{s.t. }&\theta^{*}\in\arg\min_{\theta}F(\theta,\mathcal{S})\\ &\bar{\mathcal{Q}}=\left\{\mathcal{Q}^{\prime}:I(\mathcal{Q}^{\prime},\mathcal{S}\cup\mathcal{Q}_{N})\leq r\right\}\end{array}$
@@
$|O(\theta^{*},S^{*})-\hat{O}(\hat{\theta},\hat{S})|\leq C$
where $O$ and $\hat{O}$ refer to the optimal value of the data driven DRO (5) and the approximation given by Algorithm 1 and 2.
Table 1: Top-1 test accuracy on standard testing sets. ${}^{\dagger}$ indicates the result is reported based on our runs. All the results are reported based on the average of 5 runs. ${}^{\mathcal{R}}$ indicates the proposed robust dataset distillation method applied. The better results between baseline and the proposed method are marked in bold.
Dataset IPC Random DSA DM KIP IDC ${}^{\mathcal{R}}$ GLaD ${}^{\mathcal{R}}$
SVHN 1 14.6 27.5 24.2 57.3 68.5 68.9±${}_{\pm 0.4}$ ${}^{\dagger}$ 35.7±${}_{\pm 0.3}$
10 35.1 79.2 72.0 75.0 87.5 88.1±${}_{\pm 0.3}$ ${}^{\dagger}$ 72.5±${}_{\pm 0.4}$
50 70.9 84.4 84.3 80.5 90.1 90.8±${}_{\pm 0.4}$ ${}^{\dagger}$ 76.6±${}_{\pm 0.3}$
CIFAR-10 1 14.4 28.7 26.0 49.9 ${}^{\dagger}$ 51.3±${}_{\pm 0.3}$ ${}^{\dagger}$ 29.2±${}_{\pm 0.8}$
10 37.2 52.1 53.8 62.7 67.5 68.6±${}_{\pm 0.2}$ ${}^{\dagger}$ ${}_{\pm 0.5}$
50 56.5 60.6 65.6 68.6 74.5 75.3±${}_{\pm 0.5}$ ${}^{\dagger}$ 62.5±${}_{\pm 0.7}$
ImageNet-10 1 ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ 58.2±${}_{\pm 1.2}$ ${}^{\dagger}$ 36.4±${}_{\pm 0.8}$
10 46.9 52.7 52.3 - 72.8 74.6±${}_{\pm 0.9}$ ${}^{\dagger}$ 55.2±${}_{\pm 1.1}$
@@ 3.4 Distributionally Robust Optimization and Subgroup Coverage
Let us return to (3):
$\begin{array}[]{rl}\min\limits_{\mathcal{S}}\max\limits_{\mathcal{Q}\in\mathbb{Q}}&\mathbb{E}[F(\theta^{*},\mathcal{Q})]\\ \text{s.t. }&\theta^{*}\in\arg\min_{\theta}F(\theta,\mathcal{S})\end{array}$
@@
We can directly apply Theorem 7 of  (Van Parys et al., 2021) to deduce that the result satisfies:
@@ 4 Numerical Results
Table 2: Top-1 test accuracy on robustness testing sets. “Gain” denotes the performance gain of applying our proposed method. Except for absolute values, the relative fluctuation compared with the standard case is also reported as subscripts. The experiments are conducted under the IPC setting of 10. ${}^{\mathcal{R}}$ indicates the proposed robust dataset distillation method applied. The better results between baseline and the proposed method are marked in bold.
Dataset Setting Random IDC ${}^{\mathcal{R}}$ Gain GLaD ${}^{\mathcal{R}}$ Gain
CIFAR-10 Standard 37.2 67.5 68.6 1.1 46.7 50.2 3.5
@@
ImageNet-10 Standard 46.9 72.8 74.6 1.8 50.9 55.2 4.3
@@
[…]
4.1 Implementation Details
The proposed method can be applied to various popular dataset distillation frameworks. In this paper, we perform experiments on pixel-level distillation method IDC (Kim et al., 2022) and latent-level method GLaD (Cazenavette et al., 2023) to substantiate the consistent efficacy of our method. The experiments are conducted on popular dataset distillation benchmarks, namely SVHN, CIFAR-10 and ImageNet (Yuval, 2011; Krizhevsky et al., 2009; Deng et al., 2009). The images of SVHN and CIFAR-10 are resized to 32$\times$32, while those from ImageNet-10 are resized to 128$\times$128, representing diverse resolution scenarios. During the distilling process, we generally adhere to the hyperparameter settings of the original methods. Due to the CVaR loss calculation involving an ample number of samples, the mini-batch size during model updating is increased to 256. In cases where the IPC setting is less than 10, the cluster number in Eq. 2 is set equal to IPC. For larger IPCs, the cluster number is fixed at 10, with 10 random synthesized samples chosen as the clustering centers. The ratio $\alpha$ in CVaR loss is set to 0.8. The validation protocol aligns with the configurations in (Kim et al., 2022), while the results of GLaD are all re-obtained following the same protocol.
4.2 Results on Standard Benchmark
We first evaluate our proposed method on standard testing sets, including SVHN, CIFAR-10, and ImageNet-10 in Tab. 1. The ImageNet-10 split follows the configuration outlined in IDC (Kim et al., 2022). Except for IDC and GLaD, we also compare the validation results with DSA, DM, and KIP (Zhao & Bilen, 2021, 2023; Nguyen et al., 2021). To fully demonstrate the performance of the compared methods, we implement distillation using the architectures specified in the original papers. For both IDC and GLaD on SVHN and CIFAR-10, a 3-layer ConvNet (Gidaris & Komodakis, 2018) is employed. On ImageNet, ResNet-10 (He et al., 2016) is utilized for IDC, while a 5-layer ConvNet is used for GLaD. After distillation, we conduct validation procedures on ConvNet-3 for 32$\times$32 datasets and ResNet-10 for ImageNet following the IDC protocol, ensuring a fair and consistent basis for comparison.
The incorporation of CVaR loss consistently enhances performance across all scenarios. Notably, under the IPC of 10, facilitated by the supervision from segregated sub-clusters, our proposed method demonstrates the most substantial performance improvement over baselines. In the case of 1 IPC, where only a single synthetic sample is available for sub-cluster construction, the introduction of additional CVaR loss still yields a significant performance boost, despite the absence of guidance from multiple sub-clusters. Our proposed method is especially effective on ImageNet-10, characterized by finer class divisions and greater intra-class variation compared with CIFAR-10. On ImageNet-10 we achieve an average top-1 accuracy gain of 3.1%, further elevating the state-of-the-art baseline in DD methods.
4.3 Results on Robustness Settings
@@
In addition to standard validation scenarios, a notable advantage offered by our proposed method is the robustness against various domain shifts. This property is assessed through multiple protocols. Firstly, as suggested before, we present validation results on different partitions of the testing set. The worst accuracy among the clustered testing sub-sets is reported to demonstrate the robustness of distilled data, denoted as “Cluster-min” in Tab. 2. In the sub-scripts, we report the performance drop compared with the standard case. Several key observations emerge from the experiment results. (1) In comparison to random images, the Cluster-min accuracy of baseline DD methods exhibits improvement alongside the standard performance. It suggests that by condensing the knowledge from original data into informative distilled samples, DD methods contribute to enhanced data robustness. (2) Compared with CIFAR-10, the performance gap between the standard case and the worst sub-cluster on ImageNet-10 is more pronounced. This discrepancy can be attributed to a higher incidence of ID-unrelated interruptions in ImageNet-10, resulting in larger domain-shifts between sub-clusters and the original distribution. This finding aligns with the observation in Tab. 1. (3) With our proposed robust method applied, not only is the cluster-min performance improved, but the performance drop from the standard case is also significantly mitigated compared with baselines. It suggests exceptional overall generalization and robustness conferred by our method.
@@
Table 3: Cross-architecture generalization performance comparison. The experiment is conducted on ImageNet-10 under 10 IPC settings. The distilling architecture for IDC is ResNet-10, while for GLaD is ConvNet-5. ${}^{\mathcal{R}}$ indicates the proposed robust dataset distillation method applied. The better results between baseline and the proposed method are marked in bold.
@@
IDC 71.9 72.8 70.8 55.2 64.5
${}^{\mathcal{R}}$ 72.6 74.6 72.7 56.4 65.6
GLaD 48.2 50.9 51.2 36.8 44.2
${}^{\mathcal{R}}$ 51.9 55.2 53.6 39.2 46.3
Table 4: Ablation study on CVaR loss. The experiment is conducted on CIFAR-10 and ImageNet-A under 10 IPC. The baseline model is GLaD. “CE” denotes cross entropy loss, and “CL” refers to cluster accuracy. The best results are marked in bold.
@@
CE CVaR CVaR Acc CL Acc CL
-avg -max -min -min
- - 46.7 40.2 53.9 40.5
- 49.1 45.3 56.4 43.9
- 47.9 44.2 56.1 43.7
50.2 46.7 57.5 45.8
Table 5: Validation accuracy on more ImageNet sub-sets in comparison with GLaD. The data is distilled with ConvNet-5 and evaluated with ResNet-10 with an IPC of 10. ${}^{\mathcal{R}}$ indicates the proposed robust dataset distillation method applied. The better results between baseline and the proposed method are marked in bold.
Dataset Acc Cluster-min
GLaD ${}^{\mathcal{R}}$ GLaD ${}^{\mathcal{R}}$
ImageNet-A 53.9 57.5 40.5 45.8
ImageNet-B 50.3 53.8 42.9 47.0
ImageNet-C 49.2 51.3 28.2 32.8
ImageNet-D 39.1 40.9 27.0 31.3
ImageNet-E 38.9 41.1 25.8 30.0
4.4 Cross-architecture Generalization
@@
In addition to enhanced robustness against domain-shifted data, the incorporation of CVaR loss also yields better cross-architecture generalization capabilities. We assess multiple network architectures including ConvNet-5, ResNet, ViT (Dosovitskiy et al., 2020) and VGG11 (Simonyan & Zisserman, 2014), and compare the performance with and without our proposed method in Tab. 3. Despite the distinct distilling architectures employed in IDC and GLaD, both achieve their highest accuracy on ResNet-10. Notably, the proposed robust distillation method consistently enhances performance across all architectures, showcasing remarkable cross-architecture generalization capabilities.
4.5 Ablation Study
We conduct an ablation study on the incorporation of CVaR loss in Tab. 4. In addition to CIFAR-10, we also report results on the ImageNet-A sub-set (Cazenavette et al., 2023). Both accuracy on the standard testing set and the minimum cluster accuracy are presented.

### v3

@@ 3 Analysis
=$\mathbb{Q}=\{\mathcal{Q}_{i}\},\,\text{ with }Q_{i}=\text{supp}(\mathcal{Q}_{i}),\,\bigcup_{i=1}^{q}Q_{i}=D,\,\underline{s}\leq\lambda(Q_{i})\leq\bar{s}$ (5)
@@
The population distribution $\mathcal{D}$ has bounded support, i.e., $\lambda(D)<\infty$. The loss $F$ is Lipschitz continuously differentiable with respect to the first argument (the parameters) and continuous with respect to the second argument (the input features and labels). Finally, the partitions are not probabilistically small, i.e., there exists $p_{q_{0}}$ such that $\mathbb{P}_{\mathcal{D}}[A\in\mathcal{Q}_{i}]\geq p_{q_{0}}$ for all $i$.
@@
$\lim\sup\limits_{t\to\infty}(\text{supp}(\mathcal{Q})\cap\bar{\mathcal{T}}_{t})=\text{supp}(\mathcal{Q})$ (6)
@@
$\begin{array}[]{rl}\min\limits_{\mathcal{S}}\max\limits_{\mathcal{Q}\in\mathbb{Q}}\mathbb{E}[F(\theta^{*},\mathcal{Q})]\quad\text{s.t. }\theta^{*}\in\arg\min_{\theta}F(\theta,\mathcal{S}).\end{array}$ (7)
Let us briefly consider the standard circumstance by which Algorithm 1 converges. By considering each iteration as two players’ best response to a cooperative Nash game, we can ensure convergence asymptotically of iteratively solving the two optimization problems, as given in Nash (1953).
Theorem 3.3.
Under the circumstance by which the Morse-Saard condition holds (Souček & Souček, 1972) and so the optimal set $\{\mathcal{S}^{*}(\theta)\},\{\theta^{*}(\mathcal{S})\}$ is compact (possibly finite) for all $\theta,\mathcal{S}$, then Algorithm 2 converges to a fixed point of equation 7.
A remark on the time complexity of the Algorithm. Using CVaR instead of a sample average only takes a constant multiple of operations on the subsample. Clustering itself can be done polynomial in the number of samples. Since the properties of the model and loss function, that is, nonconvex and nonsmooth, are fundamentally unchanged, there is no change to the iteration complexity.
Now we will discuss generalization and DRO. Recall that there are two layers of generalization: on the one hand, we are solving a problem robust with respect to the choice of $\mathcal{Q}\in\mathbb{Q}$, and on the other hand, we are considering the population error rather than an empirical loss.
@@
that is ultimately used to steer the synthetic dataset, as the DD task is finding a dataset on whose associated $\theta$-minimizer also minimizes this quantity. In the analysis, we shall argue about the validity of these criteria as far ensuring the generalization as well as group robustness properties of the solution of equation 7.
To begin with, we rewrite equation 7 as:
$\begin{array}[]{rl}\min\limits_{\mathcal{S}}\max\limits_{\mathcal{Q}\in\mathbb{Q}}\max_{\mathcal{Q}^{\prime}\in\bar{\mathcal{Q}}}\mathbb{E}[F(\theta^{*},\mathcal{Q}^{\prime})]\quad\text{s.t. }&\theta^{*}\in\arg\min_{\theta}F(\theta,\mathcal{S})\\ &\bar{\mathcal{Q}}=\left\{\mathcal{Q}^{\prime}:I(\mathcal{Q}^{\prime},\mathcal{S}\cup\mathcal{Q}_{N})\leq r\right\}\end{array}$ (9)
where $r>0$ is some bound and we replace the inner problem to be evaluated on the entire training dataset with a data-driven DRO. In the appendix we present several results in the literature that indicate how the DRO to an empirical risk minimization problem exhibits generalization guarantees and hence is a valid auxiliary criterion for minimizing equation 7.
@@ 3.1 Large Deviations and Solving the bilevel DRO
To justify the clustering and risk measure minimization algorithm as an appropriate procedure for solving equation 9, we apply some theoretical analysis on the relationship between DRO and Large Deviations Principles (LDP). LDPs (e.g.Deuschel & Stroock (2001)) define an exponential asymptotic decay of the measure of the tails of an empirical distribution with respect to a sample size.
@@
$\begin{array}[]{l}\lim\sup\limits_{N\to\infty}\frac{1}{N}\log\mathbb{P}_{\mathcal{Q}}^{\infty}\left(F(\theta^{*}(\mathcal{S}),\mathbb{P}_{\mathcal{Q}})>F(\theta^{*}(\mathcal{S}),\mathcal{S}\cup\mathcal{Q}_{N})\right)\leq-r,\,\forall\mathcal{Q}\in\mathbb{Q}\end{array}$ (10)
for some $r>0$. This states that for all partitions $\mathcal{Q}$ of the population space partition $\mathbb{Q}$, there is an exponential asymptotic decay of the probability of one-sided sample error (i.e., samples of size $N$ from the population $\mathbb{P}_{\mathcal{Q}}$) relative to the computed loss on the synthetic and training data points $\mathcal{Q}_{N}$, as the number of training data points grows asymptotically.
@@
Algorithm equation 2 satisfies, under Assumptions 3.1 and 3.2, the LDP equation 10.
Next, we relate Algorithm equation 2 together with 3.1 to the DRO problem equation 9.
@@ 3.2 Large Deviations and Distributionally Robust Optimization
Applying Assumption 3.2, the LDP, and the continuity of $F$ with respect to the data, Assumption 3.1 (see also the proof of Lemma 1 as well as Example 3 in Duchi & Namkoong (2021)) we can bound the quantity:
$\mathbb{P}(F(\theta^{*},q)>F(\theta^{*},\mathcal{S}\cup\mathcal{Q}_{N})$ (11)
@@
This implies that one can bound the objective value of the DRO problem of interest equation 9:
$\begin{array}[]{rl}\min\limits_{\mathcal{S}}\max\limits_{\mathcal{Q}\in\mathbb{Q}}\max_{\mathcal{Q}^{\prime}\in\bar{\mathcal{Q}}}\mathbb{E}[F(\theta^{*},\mathcal{Q}^{\prime})]\quad\text{s.t. }&\theta^{*}\in\arg\min_{\theta}F(\theta,\mathcal{S})\\ &\bar{\mathcal{Q}}=\left\{\mathcal{Q}^{\prime}:I(\mathcal{Q}^{\prime},\mathcal{S}\cup\mathcal{Q}_{N})\leq r\right\}\end{array}$ (12)
@@
$|O(\theta^{*},S^{*})-\hat{O}(\hat{\theta},\hat{S})|\leq C$ (13)
where $O$ and $\hat{O}$ refer to the optimal value of the data driven DRO equation 9 and the approximation given by Algorithm 1 and 2.
@@ 3.3 Distributionally Robust Optimization and Subgroup Coverage
Let us return to equation 7:
$\begin{array}[]{rl}\min\limits_{\mathcal{S}}\max\limits_{\mathcal{Q}\in\mathbb{Q}}\mathbb{E}[F(\theta^{*},\mathcal{Q})]\quad\text{s.t. }\theta^{*}\in\arg\min_{\theta}F(\theta,\mathcal{S})\end{array}$
@@
We can directly apply Theorem 7 in the work of  Van Parys et al. (2021) to deduce that the result satisfies:
@@ 4 Numerical Results
Table 1: Top-1 test accuracy on robustness testing sets. “Gain” denotes the performance gain of applying our proposed method. Except for absolute values, the relative fluctuation compared with the standard case is also reported as subscripts. The experiments are conducted under the IPC setting of 10. R indicates the proposed robust dataset distillation method applied. The better results between baseline and the proposed method are marked in bold.
Dataset Setting Random IDC IDCR Gain GLaD GLaDR Gain
CIFAR-10 Standard 37.2±0.8 67.5±0.5 68.6±0.2 1.1 46.7±0.6 50.2±0.5 3.5
@@
ImageNet-10 Standard 46.9±0.7 72.8±0.6 74.6±0.9 1.8 50.9±0.9 55.2±1.1 4.3
Cluster-min 31.2±1.0 61.4±1.0↓11.4 65.7±${}_{\pm 0.5\ \ \downarrow 8.9}$ 4.3↑2.5 34.9±0.9↓16.0 47.1±${}_{\pm 1.2\ \ \downarrow 8.1}$ 12.2↑7.9
@@
Table 2: Top-1 test accuracy on standard testing sets. indicates the result is reported based on our runs. R indicates the proposed robust dataset distillation method applied on the baseline. The better results between baseline and the proposed method are marked in bold.
Dataset IPC Random DSA DM KIP IDC IDCR GLaD GLaDR
SVHN 1 14.6 27.5 24.2 57.3 68.5±0.9 68.9±0.4 32.5±0.5 35.7±0.3
10 35.1 79.2 72.0 75.0 87.5±0.3 88.1±0.3 68.2±0.4 72.5±0.4
50 70.9 84.4 84.3 80.5 90.1±0.1 90.8±0.4 71.8±0.6 76.6±0.3
CIFAR-10 1 14.4 28.7 26.0 49.9 50.6±0.4 51.3±0.3 28.0±0.8 29.2±0.8
10 37.2 52.1 53.8 62.7 67.5±0.5 68.6±0.2 46.7±0.6 50.2±0.5
50 56.5 60.6 65.6 68.6 74.5±0.1 75.3±0.5 59.9±0.9 62.5±0.7
ImageNet-10 1 23.2 30.6 30.2 - 54.4±1.1 58.2±1.2 33.5±0.9 36.4±0.8
10 46.9 52.7 52.3 - 72.8±0.6 74.6±0.9 50.9±1.0 55.2±1.1
In this section we present numerical results to validate the efficacy of the proposed robust dataset distillation method. Implementation details are listed in Sec. H.
Results on Robustness Settings
@@
We first show the notable advantage offered by our proposed method that is the robustness against various domain shifts. This property is assessed through multiple protocols. Firstly, as suggested before, we present validation results on different partitions of the testing set. A clustering process is conducted to divide the original testing set into multiple sub-sets. We test the performance on each of them and report the worst accuracy among the sub-sets to demonstrate the robustness of distilled data, denoted as “Cluster-min” in Tab. 1. In the sub-scripts, the performance drop compared with the standard case is reported. Several key observations emerge from the experiment results. (1) Compared with random images, the Cluster-min accuracy of baseline DD methods exhibits improvement alongside the standard performance. It suggests that by condensing the knowledge from original data into informative distilled samples, DD methods contribute to enhanced data robustness. (2) Compared with CIFAR-10, the performance gap between the standard case and the worst sub-cluster on ImageNet-10 is more pronounced. This discrepancy can be attributed to a higher incidence of ID-unrelated interruptions in ImageNet-10, resulting in larger domain-shifts between sub-clusters and the original distribution. This finding aligns with the observation in Tab. 2. (3) With our proposed robust method applied, not only is the cluster-min performance improved, but the performance drop from the standard case is also significantly mitigated compared with baselines. It suggests exceptional overall generalization and robustness conferred by our method.
Table 3: Top-1 test accuracy on subpopulation shift benchmarks MetaShift and ImageNetBG. All the results are reported based on the average of 5 runs. R indicates the proposed robust dataset distillation method applied. The better results between baseline and the proposed method are marked in bold.
Metric MetaShift ImageNetBG
GLaD GLaDR GLaD GLaDR
Average Accuracy 58.6±2.3 62.2±1.2 41.7±1.5 45.5±1.1
Worst-group Accuracy 51.3±1.8($\downarrow 7.3$) 57.0±1.0($\downarrow 5.2$) 32.2±1.6($\downarrow 9.5$) 38.6±1.0($\downarrow 6.9$)
Subpopulation Shift Experiments
In addition to the previous standard classification tasks, we also extend the method to subpopulation shift benchmarks MetaShift and ImageNetBG (Yang et al., 2023; Liang & Zou, 2022; Xiao et al., 2021). The two benchmarks consider the spurious correlation and attribute generalization problems, respectively. We use GlaD as the baseline method to distill 50 images for each class. Subsequently, we conduct the standard evaluation as in the other datasets, with the results reported in Table 3. The proposed robust dataset distillation algorithm not only improves the average accuracy, but also yields a smaller worst-group accuracy margin compared with the baseline method. The results further suggest that RDD can enhance the robustness of distilled data for datasets with subpopulation shift issues.
@@
Table 4: (a) Cross-architecture generalization performance comparison. The experiment is conducted on ImageNet-10 under 10 IPC settings. The distilling architecture for IDC is ResNet-10, while for GLaD is ConvNet-5. R indicates the proposed robust dataset distillation method applied. (b) Ablation study on CVaR loss. The experiment is conducted on CIFAR-10 and ImageNet-A under 10 IPC. The baseline model is GLaD. “CE” denotes cross entropy loss, and “CL-min” refers to the worst sub-cluster accuracy. “aCVaR” and “mCVaR” refer to average and maximum CVaR loss, respectively. The best results are marked in bold.
@@
IDC 71.9±0.8 72.8±0.5 70.8±1.0 55.2±1.2 64.5±0.6
IDCR 72.6±0.6 74.6±0.7 72.7±0.8 56.4±1.1 65.6±0.5
GLaD 48.2±0.7 50.9±0.5 51.2±0.4 36.8±0.9 44.2±1.0
GLaDR 51.9±0.7 55.2±0.6 53.6±0.5 39.2±0.9 46.3±0.8
@@
CE aCVaR mCVaR Acc CL-min Acc CL-min
- - 46.7±0.6 40.2±0.8 53.9±0.6 40.5±0.7
- 49.1±0.7 45.3±1.0 56.4±0.5 43.9±0.8
- 47.9±0.8 44.2±0.9 56.1±0.8 43.7±0.9
50.2±0.6 46.7±0.8 57.5±0.7 45.8±0.5
Results on Standard Benchmark
We then evaluate our proposed method on standard testing sets, including SVHN (Sermanet et al., 2012), CIFAR-10 (Krizhevsky et al., 2009), ImageNet-10 (Deng et al., 2009), and Tiny-ImageNet (Deng et al., 2009) in Table 2. The ImageNet-10 split follows the configuration outlined in IDC (Kim et al., 2022). We use IDC (Kim et al., 2022) and GLaD (Cazenavette et al., 2023) as baselines in this section, representing distilling at the pixel level and the latent level, respectively. Additionally, the validation results are also compared with DSA, DM, and KIP (Zhao & Bilen, 2021; 2023b; Nguyen et al., 2021).
The incorporation of CVaR loss consistently enhances performance across all scenarios. Notably, under the IPC of 10, facilitated by the supervision from segregated sub-clusters, our proposed method demonstrates the most substantial performance improvement over baselines. In the case of 1 IPC, where only a single synthetic sample is available for sub-cluster construction, the optimization comes back to a DRO problem. And minimizing CVaR still captures tail risk and helps improve the validation performance, despite the absence of guidance from multiple sub-clusters. Our proposed method is especially effective on ImageNet-10, characterized by finer class divisions and more substantial intra-class variation compared with CIFAR-10. On ImageNet-10 we achieve an average top-1 accuracy gain of 3.1%, further elevating the state-of-the-art baseline in DD methods. More results are presented in the Appendix Section D.
Cross-architecture Generalization
@@
In addition to enhanced robustness against domain-shifted data, the incorporation of CVaR loss also yields better cross-architecture generalization capabilities. We assess multiple network architectures including ConvNet-5, ResNet, ViT (Dosovitskiy et al., 2020) and VGG11 (Simonyan & Zisserman, 2014), and compare the performance with and without our proposed robust method in Table 4(a). Despite the different distilling architectures employed in IDC and GLaD, both methods achieve their highest accuracy on ResNet-10. Notably, the proposed robust distillation method consistently enhances performance across all architectures, showcasing remarkable cross-architecture generalization capabilities.
Ablation Study